Arcadiouse (arcadiouse):

Can anyone share their ways to solving equations with variables on both sides?

OpenStudy (4everaddicted2anime):

What kind of equations are you talking about?

OpenStudy (faiqraees):

Substitution method is the best

OpenStudy (bobo-i-bo):

Depends very much! If it's a quadratic equation then you could use the quadratic formula or complete the square. Factorization is a very important and useful tool. Also, difference of two squares is quite a good thing to spot as well.

Arcadiouse (arcadiouse):

Like equations like 2x + 3 = 5x -2. Equations like that...

OpenStudy (4everaddicted2anime):

What you do is move the x's to one side and the other terms to the other side, then divide by the x's final coeficcient

OpenStudy (4everaddicted2anime):

on both sides for the final step

OpenStudy (agent0smith):

Add or subtract things from both sides, until you end up with all the variables on one side, and all the constants on the other side. That's your goal. Get variables on one side, constants on the other.

OpenStudy (anonymous):

Sometimes, the unknown quantity will appear on both sides of an equation. This is where the properties learned in 5.1 and 5.2 come in handy. A quantity with a variable can be treated just like a quantity without variables -- a quantity with a variable follows all the rules learned in the last two sections. For example, we can add a quantity with a variable to both sides without changing the equation or the values that make it true: 15 - x = 4x 15 - x + x = 4x + x 15 + 0x = 5x 15 = 5x 3 = x x = 3